# Toward Global Convergence of Gradient EM for Over-Parameterized Gaussian Mixture Models

### v1

@@ B.1 Proofs for global convergence analysis
For any $\|x\|\leq 2\sqrt{d}$, we have $\exp(-\|x-\mu_{1}\|^{2}/2)\geq\exp(-(\|x\|+\|\mu_{1}\|)^{2}/2)\geq\exp(-9d/2)$, while for $\forall i\neq 1$, $\exp(-\|x-\mu_{i}\|^{2}/2)\leq\exp(-(\|\mu_{i}\|-\|x\|)^{2}/2)\leq\exp(-(10\sqrt{d}-2\sqrt{d})^{2}/2)=\exp(-32d)$. Since $\psi_{i}(x)\propto\exp(-\|x-\mu_{i}\|^{2}/2)$ we have
$\|x\|\leq 2\sqrt{d}\Rightarrow\psi_{i}(x)\leq\frac{\exp(-\|x-\mu_{i}\|^{2}/2)}{\exp(-\|x-\mu_{1}\|^{2}/2)}\leq\frac{\exp(-32d)}{\exp(-9d/2)}\leq\exp(-25d),\forall i\neq 1.$
Therefore the first term in (36) can be bounded as $\mathbf{E}_{x}\left[\sum_{k\in[n]}\psi_{k}(x)\|\mu_{k}\|\Bigg{|}\|x\|\leq 2\sqrt{d}\right]\leq\|\mu_{1}\|+\exp(-25d)\sum_{i\neq 1}\|\mu_{i}\|.$
On the other hand, by tail bound of the norm of Gaussian vectors (see Lemma 8 of [Yan et al., ]) we have $\Pr\left[\|x\|>2\sqrt{d}\right]\leq\exp(-d)$. Putting everything together, (39) can be further bounded as
$\begin{split}\|\nabla_{\mu_{i}}\mathcal{L}(\bm{\mu})\|\leq\|\mu_{1}\|+\exp(-25d)\sum_{i\neq 1}\|\mu_{i}\|+\exp(-d)\sum_{i\in[n]}\|\mu_{i}\|\leq 2\|\mu_{1}\|+2\exp(-d)\sum_{i\neq 1}\|\mu_{i}\|.\end{split}$
@@
For any $n=2l+1$, consider gradient EM initialized at point $\mu_{1}(0)=0,\mu_{2}(0)=\cdots=\mu_{l+1}(0)=12\sqrt{d}e_{1},\mu_{l+2}(0)=\cdots=\mu_{2l+1}(0)=-12\sqrt{d}e_{1}$, where $e_{1}=(1,0,\ldots,0)^{\top}$ is a standard unit vector. Then population gradient EM will be trapped in a bad local region around $\bm{\mu}(0)$ for exponentially long time $T=\frac{1}{15n\eta}e^{d}=\frac{1}{15n\eta}\exp(\Theta(\mu_{\max}^{2}(0)))$. More rigorously, for any $0\leq t\leq T$, we have
$\|\mu_{i}(t)\|\geq 10\sqrt{d},\;\;\forall\;i\neq 1.$
@@
$\mu_{1}(t)=0,\mu_{2}(t)=\ldots=\mu_{l+1}(t)=-\mu_{l+2}(t)=\ldots=-\mu_{2l+1}(t)$ (40)
$\forall i,\;\|\mu_{i}(t)-\mu_{i}(0)\|\leq\eta t(30\sqrt{d}ne^{-d}).$ (41)
@@
Proof of (40). Due to the induction hypothesis, one can see from direct calculation that $\forall x,\psi_{1}(x|\bm{\mu}(t))=\psi_{1}(-x|\bm{\mu}(t)),\psi_{2}(x|\bm{\mu}(t))=\cdots=\psi_{l+1}(x|\bm{\mu}(t))=\psi_{l+2}(-x|\bm{\mu}(t))=\cdots=\psi_{2l+1}(-x|\bm{\mu}(t))$. For the ease of notation, we denote $\psi_{+}(x)\coloneqq\psi_{2}(x|\bm{\mu}(t))=\cdots=\psi_{l+1}(x|\bm{\mu}(t)),\psi_{-}(x)\coloneqq\psi_{l+2}(x|\bm{\mu}(t))=\cdots=\psi_{2l+1}(x|\bm{\mu}(t)),\mu^{+}\coloneqq\mu_{2}(t)=\ldots=\mu_{l+1}(t),\mu^{-}\coloneqq\mu_{l+2}(t)=\ldots=-\mu_{2l+1}(t).$ Then $\mu^{+}=-\mu^{-},\psi_{+}(x)=\psi_{-}(-x)$. Consequently
$\begin{split}&\quad\nabla_{\mu_{1}}\mathcal{L}(\bm{\mu}(t))=\mathbf{E}_{x}\left[\psi_{1}(x|\bm{\mu}(t))\sum_{k\in[n]}\psi_{k}(x|\bm{\mu}(t))\mu_{k}(t)\right]=\mathbf{E}_{x}\left[\psi_{1}(x)(l\psi_{+}(x)\mu^{+}+l\psi_{-}(x)\mu^{-})\right]\\ &=\frac{1}{2}\mathbf{E}_{x}\left[\psi_{1}(x)(l\psi_{+}(x)\mu^{+}+l\psi_{-}(x)\mu^{-})+\psi_{1}(-x)(l\psi_{+}(-x)\mu^{+}+l\psi_{-}(-x)\mu^{-})\right]\\ &=\frac{1}{2}\mathbf{E}_{x}\left[\psi_{1}(x)(l\psi_{+}(x)(\mu^{+}+\mu^{-})+l\psi_{-}(x)(\mu^{+}+\mu^{-}))\right]=0\Rightarrow\mu_{1}(t+1)=\mu_{1}(t)=0.\end{split}$
Similarly, $\forall 2\leq i\leq l+1,l+2\leq j\leq 2l+1$ we have
$\begin{split}&\quad\nabla_{\mu_{i}}\mathcal{L}(\bm{\mu}(t))=\mathbf{E}_{x}\left[\psi_{i}(x|\bm{\mu}(t))\sum_{k\in[n]}\psi_{k}(x|\bm{\mu}(t))\mu_{k}(t)\right]=\mathbf{E}_{x}\left[\psi_{+}(x)(l\psi_{+}(x)\mu^{+}+l\psi_{-}(x)\mu^{-})\right]\\ &=\mathbf{E}_{x}\left[\psi_{+}(-x)(l\psi_{+}(-x)\mu^{+}+l\psi_{-}(-x)\mu^{-})\right]=-\mathbf{E}_{x}\left[\psi_{-}(x)(l\psi_{+}(x)\mu^{+}+l\psi_{-}(x)\mu^{-})\right]=-\nabla_{\mu_{j}}\mathcal{L}(\bm{\mu}(t)).\end{split}$
This indicates that $\nabla_{\mu_{2}}\mathcal{L}(\bm{\mu}(t))=\cdots=\nabla_{\mu_{l+1}}\mathcal{L}(\bm{\mu}(t))=-\nabla_{\mu_{l+2}}\mathcal{L}(\bm{\mu}(t))=\cdots=-\nabla_{\mu_{2l+1}}\mathcal{L}(\bm{\mu}(t))$, combined with the induction hypothesis we have $\mu_{2}(t+1)=\ldots=\mu_{l+1}(t+1)=-\mu_{l+2}(t+1)=\ldots=-\mu_{2l+1}(t+1)$, (40) is proved.
@@
By induction hypothesis, we have $\forall i,\;\|\mu_{i}(t)-\mu_{i}(0)\|\leq\eta t\cdot(30\sqrt{d}ne^{-d})\leq\eta T\cdot(30\sqrt{d}ne^{-d})\leq 2\sqrt{d}.$ So $\forall i\neq 1,\|\mu_{i}(t)\|\leq\|\mu_{i}(0)\|+2\sqrt{d}<15\sqrt{d}$. Then by Lemma 16, $\forall i\in[n]$ we have
$\|\nabla_{\mu_{i}}\mathcal{L}(\bm{\mu}(t))\|\leq 2\|\mu_{1}(t)\|+2\exp(-d)\sum_{i\neq 1}\|\mu_{i}(t)\|\leq 2n\exp(-d)\cdot 15\sqrt{d}=30\sqrt{d}ne^{-d},$
note that here we used $\mu_{1}(t)=0$. Therefore by the induction hypothesis we have $\|\mu_{i}(t+1)-\mu_{i}(0)\|\leq\eta t\cdot(30\sqrt{d}ne^{-d})+\eta\|\nabla_{\mu_{i}}\mathcal{L}(\bm{\mu}(t))\|\leq\eta(t+1)\cdot(30\sqrt{d}ne^{-d})$, (41) is proven.
By (41), $\forall i\neq 1,0\leq t\leq T$ we have $\|\mu_{i}(t)\|\geq\|\mu_{i}(0)\|-\|\mu_{i}(t)-\mu_{i}(0)\|\geq 12\sqrt{d}-\eta T(30\sqrt{d}ne^{-d})\geq 12\sqrt{d}-2\sqrt{d}=10\sqrt{d}.$ Our proof is done. ∎

### v2

@@ B.1 Proofs for global convergence analysis
For any $\|x\|\leq 2\sqrt{d}$ and $i>2$, we have $\exp(-\|x-\mu_{i}\|^{2}/2)\geq\exp(-(\|x\|+\|\mu_{i}\|)^{2}/2)\geq\exp(-9d/2)$, while for $i\in\{1,2\}$, $\exp(-\|x-\mu_{i}\|^{2}/2)\leq\exp(-(\|\mu_{i}\|-\|x\|)^{2}/2)\leq\exp(-(10\sqrt{d}-2\sqrt{d})^{2}/2)=\exp(-32d)$. Since $\psi_{i}(x)\propto\exp(-\|x-\mu_{i}\|^{2}/2)$ we have
$\|x\|\leq 2\sqrt{d}\Rightarrow\psi_{i}(x)\leq\frac{\exp(-\|x-\mu_{i}\|^{2}/2)}{\exp(-\|x-\mu_{1}\|^{2}/2)}\leq\frac{\exp(-32d)}{\exp(-9d/2)}\leq\exp(-25d),\forall i\in\{1,2\}.$
Therefore the first term in (36) can be bounded as $\mathbf{E}_{x}\left[\sum_{k\in[n]}\psi_{k}(x)\|\mu_{k}\|\Bigg{|}\|x\|\leq 2\sqrt{d}\right]\leq(\|\mu_{3}\|+\cdots+\|\mu_{n}\|)+\exp(-25d)(\|\mu_{1}\|+\|\mu_{2}\|).$
On the other hand, by tail bound of the norm of Gaussian vectors (see Lemma 8 of [Yan et al., 2017]) we have $\Pr\left[\|x\|>2\sqrt{d}\right]\leq\exp(-d)$. Putting everything together, (39) can be further bounded as
$\begin{split}\|\nabla_{\mu_{i}}\mathcal{L}(\bm{\mu})\|&\leq(\|\mu_{3}\|+\cdots+\|\mu_{n}\|)+\exp(-25d)(\|\mu_{1}\|+\|\mu_{2}\|)+\exp(-d)\sum_{i\in[n]}\|\mu_{i}\|\\ &\leq 2(\|\mu_{3}\|+\cdots+\|\mu_{n}\|)+2\exp(-d)(\|\mu_{1}\|+\|\mu_{2}\|).\end{split}$
@@
For any $n\geq 3$, define $\tilde{\bm{\mu}}(0)=({\mu}^{\top}_{1}(0),\ldots,{\mu}^{\top}_{n}(0))$ as follows: ${\mu}_{1}(0)=12\sqrt{d}e_{1},{\mu}_{2}(0)=-12\sqrt{d}e_{1},\mu_{3}(0)=\cdots=\mu_{n}(0)=0$, where $e_{1}$ is a standard unit vector. Then population gradient EM initialized with means $\tilde{\bm{\mu}}(0)$ and equal weights $\pi_{1}=\ldots=\pi_{n}=1/n$ will be trapped in a bad local region around $\tilde{\bm{\mu}}(0)$ for exponentially long time $T=\frac{1}{30\eta}e^{d}=\frac{1}{30\eta}\exp(\Theta(U(0)))$. More rigorously, for any $0\leq t\leq T,\exists i\in[n]$ such that
$\|\mu_{i}(t)\|\geq 10\sqrt{d},\;\;$
@@
${\mu}_{1}(t)+{\mu}_{2}(t)=0,\mu_{3}(t)=\cdots=\mu_{n}(t)=0$ (40)
$\forall i,\;\|\mu_{i}(t)-\mu_{i}(0)\|\leq\eta t(60\sqrt{d}e^{-d}).$ (41)
@@
Proof of (40). Due to the induction hypothesis, one can see from direct calculation that $\forall x,$ we have $\psi_{i}(x|\bm{\mu}(t))=\psi_{i}(-x|\bm{\mu}(t))$ for $i=3,\ldots,n$, and $\psi_{1}(x|\bm{\mu}(t))=\psi_{2}(-x|\bm{\mu}(t))$.
Consequently for $\forall i>2$ we have
$\begin{split}&\quad\nabla_{\mu_{i}}\mathcal{L}(\bm{\mu}(t))=\mathbf{E}_{x}\left[\psi_{i}(x|\bm{\mu}(t))\sum_{k\in[n]}\psi_{k}(x|\bm{\mu}(t))\mu_{k}(t)\right]=\mathbf{E}_{x}\left[\psi_{i}(x)(\psi_{1}(x)\mu_{1}(t)+\psi_{2}(x)\mu_{2}(t))\right]\\ &=\frac{1}{2}\mathbf{E}_{x}\left[\psi_{i}(x)(\psi_{1}(x)\mu_{1}(t)+\psi_{2}(x)\mu_{2}(t))+\psi_{i}(-x)(\psi_{1}(-x)\mu_{1}(t)+\psi_{2}(-x)\mu_{2}(t))\right]\\ &=\frac{1}{2}\mathbf{E}_{x}\left[\psi_{i}(x)(\psi_{1}(x)(\mu_{1}(t)+\mu_{2}(t))+\psi_{2}(x)(\mu_{2}(t)+\mu_{1}(t)))\right]=0\Rightarrow\mu_{1}it+1)=\mu_{i}(t)=0.\end{split}$
Similarly, for $\mu_{1},\mu_{2}$ we have
$\begin{split}&\quad\nabla_{\mu_{1}}\mathcal{L}(\bm{\mu}(t))=\mathbf{E}_{x}\left[\psi_{1}(x|\bm{\mu}(t))\sum_{k\in[n]}\psi_{k}(x|\bm{\mu}(t))\mu_{k}(t)\right]=\mathbf{E}_{x}\left[\psi_{1}(x)(\psi_{1}(x)\mu_{1}+\psi_{2}(x)\mu_{2})\right]\\ &=\mathbf{E}_{x}\left[\psi_{2}(-x)(\psi_{2}(-x)\mu_{1}+\psi_{1}(-x)\mu_{2})\right]=-\mathbf{E}_{x}\left[\psi_{2}(-x)(\psi_{2}(-x)\mu_{2}+\psi_{1}(-x)\mu_{1})\right]=-\nabla_{\mu_{2}}\mathcal{L}(\bm{\mu}(t)).\end{split}$
This combined with the induction hypothesis implies $\mu_{2}(t+1)=-\mu_{1}(t+1)$, (40) is proved.
@@
By induction hypothesis, we have $\forall i,\;\|\mu_{i}(t)-\mu_{i}(0)\|\leq\eta t\cdot(60\sqrt{d}e^{-d})\leq\eta T\cdot(60\sqrt{d}e^{-d})\leq 2\sqrt{d}.$ So $\forall i\in\{1,2\},\|\mu_{i}(t)\|\leq\|\mu_{i}(0)\|+2\sqrt{d}<15\sqrt{d}$. Then by Lemma 15, $\forall i\in[n]$ we have
$\|\nabla_{\mu_{i}}\mathcal{L}(\bm{\mu}(t))\|\leq 2(\|\mu_{3}\|+\cdots+\|\mu_{n}\|)+2\exp(-d)(\|\mu_{1}\|+\|\mu_{2}\|)\leq 4\exp(-d)\cdot 15\sqrt{d}=60\sqrt{d}e^{-d},$
note that here we used $\mu_{3}(t)=\cdots=\mu_{n}(t)=0$. Therefore by the induction hypothesis we have $\|\mu_{i}(t+1)-\mu_{i}(0)\|\leq\eta t\cdot(60\sqrt{d}e^{-d})+\eta\|\nabla_{\mu_{i}}\mathcal{L}(\bm{\mu}(t))\|\leq\eta(t+1)\cdot(60\sqrt{d}e^{-d})$, (41) is proven.
By (41), $\forall 0\leq t\leq T$, for $i=1,2$ we have $\|\mu_{i}(t)\|\geq\|\mu_{i}(0)\|-\|\mu_{i}(t)-\mu_{i}(0)\|\geq 12\sqrt{d}-\eta T(60\sqrt{d}e^{-d})\geq 12\sqrt{d}-2\sqrt{d}=10\sqrt{d}.$ Our proof is done. ∎